# PSGait: Multimodal Gait Recognition using Parsing Skeleton

### v1

@@ Abstract
Gait recognition has emerged as a robust biometric modality due to its non-intrusive nature and resilience to occlusion. Conventional gait recognition methods typically rely on silhouettes or skeletons. Despite their success in gait recognition for controlled laboratory environments, they usually fail in real-world scenarios due to their limited information entropy for gait representations. To achieve accurate gait recognition in the wild, we propose a novel gait representation, named Parsing Skeleton. This representation innovatively introduces the skeleton-guided human parsing method to capture fine-grained body dynamics, so they have much higher information entropy to encode the shapes and dynamics of fine-grained human parts during walking. Moreover, to effectively explore the capability of the parsing skeleton representation, we propose a novel parsing skeleton-based gait recognition framework, named PSGait, which takes parsing skeletons and silhouettes as input. By fusing these two modalities, the resulting image sequences are fed into gait recognition models for enhanced individual differentiation. We conduct comprehensive benchmarks on various datasets to evaluate our model. PSGait outperforms existing state-of-the-art multimodal methods. Furthermore, as a plug-and-play method, PSGait leads to a maximum improvement of 10.9% in Rank-1 accuracy across various gait recognition models. These results demonstrate the effectiveness and versatility of parsing skeletons for gait recognition in the wild, establishing PSGait as a new state-of-the-art approach for multimodal gait recognition.
Index Terms:
@@
Gait Recognition, Multimodal Fusion, Gait Representation, Human Parsing
I Introduction
Gait recognition aims to identify the target person by the motion pattern of human walking. Compared with face and fingerprint, gait has the advantages of being remotely accessible, non-contacting, and hard to disguise, which makes it a promising biometric for applications like ID authentication and social security[1, 2].
Existing gait representations for gait recognition primarily rely on silhouette sequences [3], Gait Energy Images (GEIs)[4], 2D skeletons[5], 3D skeletons[6], etc. Although these representations perform well in laboratory environments, they often fail in unconstrained situations, i.e., gait recognition in the wild [7, 8]. This failure is attributed to the low information entropy of binary silhouettes or 2D/3D skeletons, which are insufficient for effectively encoding the shapes and dynamics of the human body during walking. Recent research has increasingly focused on multimodal fusion methods in gait recognition[9, 10, 8, 11, 12], combining silhouettes and skeletons to create a more comprehensive representation of human gait. For example, Skeletongait++[9] utilized the Skeleton Map to represent the skeleton and integrated it with the silhouette using an attention mechanism. Although this approach has shown improvements, it relied on Gaussian points to represent the skeleton, neglecting detailed information about the body parts between these points. In gait recognition, the motion dynamics of different body parts vary significantly, and this disparity provides valuable complementary information for distinguishing individuals. Therefore, it is crucial to explore new representations with higher information entropy for gait recognition in the wild. Recent studies have demonstrated that human parsing provides a fine-grained and more comprehensive representation of gait, significantly enhancing the performance of gait recognition models[13, 14, 15, 16]. Although human parsing has shown promising results in gait recognition, conventional methods typically rely on RGB images, which are highly sensitive to factors like clothing variations and environmental occlusion, thereby reducing the accuracy of gait recognition in the wild.
@@ I Introduction
Figure 1: The comparison of various gait representations. Only a single frame is displayed for brevity. Result shows the average Rank-1 on CCPG[17] for different representation inputs to GaitBase[18].
To address these issues, we propose a novel gait representation named Parsing Skeleton, which can be easily extracted through image processing methods based on the skeleton. Compared to existing gait recognition representations, the parsing skeleton offers three key advantages: 1) it contains higher information entropy at the pixel level, encoding more details about human motion; 2) it represents part-level knowledge, enabling models to learn fine-grained dynamics of different body parts during walking; 3) being derived from skeleton data, it improves the robustness and accuracy of gait recognition in real-world scenarios, outperforming existing RGB-based parsing methods. To effectively explore the capability of the parsing skeleton, we propose a parsing skeleton-based gait recognition framework named PSGait, which takes parsing skeletons and silhouettes as input. By fusing these two modalities through overlaying, the resulting image sequences are fed into gait recognition models for individual differentiation. We conduct comprehensive benchmarks on various datasets to evaluate our method against existing gait recognition approaches. PSGait achieves a maximum improvement of 5.2% in Rank-1 accuracy compared to the state-of-the-art multimodal method. Furthermore, replacing silhouettes with the fusion of silhouettes and parsing skeletons maximally leads to a 10.9% improvement in Rank-1 accuracy across various gait recognition models, underscoring the significant contribution of the plug-and-play parsing skeleton to gait recognition in real-world scenarios. These results validate the effectiveness and versatility of the parsing skeleton, establishing PSGait as a new state-of-the-art approach for multimodal gait recognition.
Figure 2: Overview of the PSGait framework. During data preprocessing, the silhouettes and skeletons are generated from video frames. In multimodal fusion, after the parsing section, the parsing skeletons are fused with silhouettes. The fused data is then fed into the gait recognition model for individual differentiation. The parsing section refines body part representations through selected points, creating circular and linear representations to enhance the robustness and accuracy of gait recognition in the wild. Both the data preprocessing and gait recognition models are interchangeable.
@@
  • To the best of our knowledge, this is the first work to explore a simple yet effective gait representation for gait recognition, namely Parsing Skeleton, which demonstrates higher information entropy compared to existing silhouette- or skeleton-based representations.
@@
  • Extensive experiments across various datasets validate the effectiveness and versatility of the parsing skeleton, establishing PSGait as a new state-of-the-art method for multimodal gait recognition.
II Related Work
@@
Multimodal gait recognition aims to fuse multiple modalities to create a more comprehensive representation of gait. Specifically, SMPLGait[8] leveraged 3D geometric information from the SMPL model to enhance the gait appearance feature learning. BiFusion[10] integrated silhouettes and skeletons, using a graph convolutional network (GCN) with a multi-scale gait graph (MSGG) to extract complementary features from both modalities. Skeletongait++[9] employed the Skeleton Map to represent the skeleton and integrated it with the silhouette using an attention mechanism.
@@
Human parsing in gait recognition typically relies on representations that contain fine-grained semantic information about individual body parts. These methods achieve simple but efficient human body parsing based on RGB images. Specifically, ParsingGait[13] combined CNNs and GCNs to effectively capture both global and local gait features, thereby enhancing gait recognition performance in real-world scenarios. GaitParsing[14] decomposed gait into specific body parts using human semantic parsing and employed a dual-branch feature extraction network to process both holistic and part-level features. LandmarkGait[16] introduced an unsupervised method that converts silhouettes into consistent landmarks for fine-grained parsing, utilizing a multi-scale feature extractor to capture both global and part-level features for gait recognition.
III Method
@@
III-A Parsing Section
@@
$\text{line}((x_{1},y_{1}),(x_{2},y_{2}),w,c)=\left\{(u,v)\mid d\leq\frac{w}{2}\right\}$ (4)
@@ III-B Fusion Strategy
$F_{\text{overlay}}(x,y)=\max\left(M(x,y),S(x,y)\right),$ (5)
@@
$F_{\text{multi}}(x,y,c)=\begin{cases}S_{c}(x,y),&\text{if }c=1,2,\dots,11\\ M(x,y),&\text{if }c=12\end{cases}$ (6)
@@
III-C Align Center
After the data fusion stage, the center alignment operation is performed to ensure that the human body contour and the resolved skeleton are centered in the image under the premise of uniform input image size. This alignment facilitates subsequent model processing and feature extraction. Given a binary silhouette image $M(x,y)$, where $M(x,y)\in\{0,1\}$, the upper bound $y_{\text{top}}$ and lower bound $y_{\text{btm}}$ of the human body region are calculated as follows:
$y_{\text{top}}=\min\left\{y\,\middle|\,\sum_{x}M(x,y)>0\right\}.$ (7)
$y_{\text{btm}}=\max\left\{y\,\middle|\,\sum_{x}M(x,y)>0\right\}.$ (8)
Next, the image is cropped based on the upper and lower bounds of the human region and resized to the target height while preserving the aspect ratio, resulting in the image $M^{\prime}(x,y)$. The center of the horizontal axis is calculated as follows:
$x_{c}=\frac{\sum_{x,y}x\cdot M^{\prime}(x,y)}{\sum_{x,y}M^{\prime}(x,y)}$ (9)
Using the horizontal axis center $x_{c}$, the image is cropped to a target size of $H^{\prime}\times W^{\prime}$.
TABLE I: Quantitative comparison of gait recognition methods on the SUSTech1K[19] is presented. The best performances are indicated in bold, and the second-best methods are underlined. †We modified Skeletongait++ to match the settings used in GaitBase.
Input Method Probe Sequence(R-1) Overall
Skeleton GaitGraph2(CVPRW2022)[20] 22.2 18.2 6.8 18.6 13.4 19.2 27.3 16.4 18.6 40.2
GaitTR(ES2023)[21] 33.3 31.5 21.0 30.4 22.7 34.6 44.9 23.5 30.8 56.0
Silhouette GaitSet(AAAI2019)[22] 69.1 68.2 37.4 65.0 63.1 61.0 67.2 23.0 65.0 84.8
GaitPart(CVPR2019)[23] 62.2 62.8 33.1 59.5 57.2 54.8 57.2 21.7 59.2 80.8
GaitGL(ICCV2021)[24] 67.1 66.2 35.9 63.3 61.6 58.1 66.6 17.9 63.1 82.8
GaitBase(CVPR2023)[18] 81.5 77.5 49.6 75.8 75.5 76.7 81.4 25.9 76.1 89.4
DeepGaitV2(Arxiv2023)[25] 83.5 79.5 46.3 76.8 79.1 78.5 81.1 27.3 77.4 90.2
[…]
Skeleton+Silhouette BiFusion(MTA2023)[10] 69.8 62.3 45.4 60.9 54.3 63.5 77.8 33.7 62.1 83.4
SkeletonGait++†(AAAI2024)[9] 84.7 82.4 46.7 81.8 81.2 82.6 85.1 44.3 81.2 93.2
[…]
TABLE II: Quantitative comparison of gait recognition methods is conducted across two authoritative datasets, CCPG[17] and Gait3D[8]. †We modified Skeletongait++ to match the settings used in GaitBase.
Input Method Testing Datasets

### v2

# PSGait: Gait Recognition using Parsing Skeleton

@@ Abstract.
Gait recognition has emerged as a robust biometric modality due to its non-intrusive nature and resilience to occlusion. Conventional gait recognition methods typically rely on silhouettes or skeletons. Despite their success in gait recognition for controlled laboratory environments, they usually fail in real-world scenarios due to their limited information entropy for gait representations. To achieve accurate gait recognition in the wild, we propose a novel gait representation, named Parsing Skeleton. This representation innovatively introduces the skeleton-guided human parsing method to capture fine-grained body dynamics, so they have much higher information entropy to encode the shapes and dynamics of fine-grained human parts during walking. Moreover, to effectively explore the capability of the Parsing Skeleton representation, we propose a novel Parsing Skeleton-based gait recognition framework, named PSGait, which takes Parsing Skeletons and silhouettes as input. By fusing these two modalities, the resulting image sequences are fed into gait recognition models for enhanced individual differentiation. We conduct comprehensive benchmarks on various datasets to evaluate our model. PSGait outperforms existing state-of-the-art multimodal methods that utilize both skeleton and silhouette inputs while significantly reducing computational resources. Furthermore, as a plug-and-play method, PSGait leads to a maximum improvement of 10.9% in Rank-1 accuracy across various gait recognition models. These results demonstrate that Parsing Skeleton offers a lightweight, effective, and highly generalizable representation for gait recognition in the wild.
@@
Gait Recognition, Gait Representation, Biometric Modality, Parsing Skeleton, Human Parsing, Multimodal Gait Recognition
copyright: acmlicensedjournalyear: 2025conference: Proceedings of the 33st ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Ireland.ccs: Computing methodologies Biometrics
1. INTRODUCTION
@@
Figure 1. The comparison of various gait representations. Only a single frame is displayed for brevity. Result shows the average Rank-1 on CCPG(Li et al., 2023b) for different representation inputs to GaitBase(Fan et al., 2023b).
Gait recognition aims to identify individuals based on their walking patterns. Compared to other biometric modalities such as face and fingerprint, gait is non-intrusive, contactless, and can be captured at a distance without subject cooperation. These advantages make gait particularly appealing for real-world applications, including person identification, access control, and public security surveillance in unconstrained environments (Niyogi and Adelson, 1994; Wan et al., 2018; Wang et al., 2003; Wu et al., 2016; Sarkar et al., 2005; Nixon et al., 1996, 2010).
Existing gait recognition methods typically rely on representations such as silhouette sequences (Yu et al., 2006), Gait Energy Images (GEIs) (Han and Bhanu, 2005), 2D skeletons (Teepe et al., 2021), or 3D skeletons (Liao et al., 2020). Among them, silhouette-based approaches are widely adopted due to their simplicity and effectiveness in controlled environments. However, as binary images, silhouettes offer limited information entropy—capturing only coarse body contours while discarding fine-grained motion and structural details. They are also highly sensitive to clothing variations, occlusion, and background noise, which often results in incomplete or distorted inputs. Furthermore, silhouette quality heavily relies on background subtraction algorithms, making them vulnerable to domain shifts and environmental inconsistencies across datasets. The lack of semantic body part information further restricts the model’s ability to learn localized motion patterns, ultimately limiting the robustness and discriminability of silhouette-based representations in the wild(Zheng et al., 2023).
Skeleton-based representations provide an alternative that abstracts human motion through joint coordinates. Recent studies have explored skeleton-based gait recognition using Graph Convolutional Networks (GCNs)(Teepe et al., 2021, 2022; Zhang et al., 2023; Fu et al., 2023), but performance still lags behind image-based methods. We argue this gap stems from the inherently low-dimensional and sparse nature of coordinate-based skeletons (e.g., 17 joints in COCO format), which results in low information entropy. Such sparse representations lack the spatial density and structural detail needed to model complex body dynamics during walking. Additionally, GCN-based methods are often constrained by limited receptive fields and struggle to capture long-range spatial dependencies, making it challenging to extract both global posture and local motion cues effectively.
To mitigate the limitations of single-modality representations, recent works have increasingly focused on multimodal fusion strategies that integrate silhouettes and skeletons (Fan et al., 2024; Peng et al., 2024; Zheng et al., 2022; Cui and Kang, 2023; Li et al., 2023a). For instance, SkeletonGait++ (Fan et al., 2024) proposed a Skeleton Map to visually encode joint coordinates and fused it with silhouettes via attention mechanisms. While this approach improves representation capacity, it still uses Gaussian points to encode joints and overlooks the motion dynamics of body regions between these points. Given that different body parts exhibit distinct motion trajectories, capturing part-level movement is crucial for discriminative gait recognition. This motivates the need for richer representations with higher information entropy that can better characterize human motion in unconstrained settings.
Human parsing has recently been explored as a promising avenue to enhance gait representation through fine-grained semantic segmentation (Zheng et al., 2023; Wang et al., 2023a; Zheng et al., 2024; Wang et al., 2023b). Parsing-based methods can provide part-level spatial detail and capture diverse motion patterns across different body regions. However, most existing approaches rely on RGB images, which are inherently sensitive to illumination changes, clothing variations, and occlusion—factors that are prevalent in real-world scenarios. This undermines the reliability and robustness of parsing-based representations in the wild.
To address these challenges, we propose a novel gait representation called Parsing Skeleton, which introduces skeleton-guided human parsing to generate dense, fine-grained body-part maps. Parsing Skeleton offers three key advantages over existing representations: (1) it exhibits higher pixel-level information entropy by encoding detailed motion cues; (2) it explicitly captures part-level semantic knowledge, enabling the model to distinguish diverse motion dynamics of different body regions; and (3) it inherits the robustness of skeletons to environmental noise and domain shifts, making it more reliable in real-world scenarios than RGB-based parsing methods. Unlike sparse coordinate-based skeletons, the Parsing Skeleton transforms joint positions into structured visual patterns, representing body parts (e.g., limbs and head) as colored line segments and circles. This enhances spatial density, preserving both local motion cues and global context. Its image-based nature makes it highly compatible with CNNs, which excel at learning spatially invariant features. Additionally, the Parsing Skeleton encodes structural priors and can be easily fused with silhouettes. In contrast, GCNs struggle with long-range dependencies and lack spatial expressiveness, making CNNs a more effective backbone.
To explore the potential of the Parsing Skeleton, we propose PSGait, a gait recognition framework that integrates Parsing Skeletons and silhouettes. By fusing these two modalities, the resulting image sequences are fed into gait recognition models for individual differentiation. We conduct comprehensive benchmarks on several datasets, demonstrating that PSGait achieves a maximum improvement of 5.2% in Rank-1 accuracy compared to the state-of-the-art multimodal methods that utilize both skeleton and silhouette inputs while significantly reducing computational resources. Moreover, replacing silhouettes with the fusion of silhouettes and Parsing Skeletons maximally leads to a 10.9% improvement in Rank-1 accuracy across various models, underscoring the significant contribution of the Parsing Skeleton to gait recognition in real-world scenarios.
@@
  • We propose Parsing Skeleton, a novel and lightweight gait representation that offers higher information entropy and fine-grained body-part awareness compared to traditional silhouette- and skeleton-based representations.
@@
  • Extensive experiments on multiple datasets validate the lightweight, effectiveness and generalization nature of Parsing Skeleton. PSGait achieves state-of-the-art performance among multimodal gait recognition methods while maintaining high computational efficiency.
2. RELATED WORK
In this section, we first introduce the existing gait representations. Then, the gait recognition methods are briefly surveyed. Finally, we summarize the human parsing methods in gait recognition.
2.1. Gait Representations
Gait recognition is a rapidly evolving field, with a primary focus on extracting distinctive gait features from visual inputs. The majority of gait representations are derived from traditional RGB images, including raw RGB images, binary silhouettes, optical flow images, 2D/3D skeletons, and human meshes. These representations are often susceptible to extraneous noise, stemming from factors such as background clutter, clothing variations, and external accessories (e.g., bags). To mitigate such influences, preprocessing stages such as background subtraction, silhouette extraction, and normalization are typically employed. Alternatively, end-to-end learning approaches, which directly learn robust gait features from raw data, are also utilized to overcome such challenges. Beyond RGB images, recent studies have explored alternative sensor modalities, such as LiDAR (Shen et al., 2023) and event cameras (Wang et al., 2022), which capture temporal and spatial information with high precision. However, the integration of these sensors remains limited due to their relatively low adoption in existing CCTV systems, thereby rendering them unsuitable for large-scale deployment in surveillance applications at present.
2.2. Gait Recognition Methods
According to the classical taxonomy, gait recognition methods can be broadly classified into two categories: model-based and appearance-based methods.
Model-based gait recognition methods utilize the underlying structure of the human body as input, such as the estimated 2D / 3D skeleton and human mesh.(Wu et al., 2025; Ariyanto and Nixon, 2011; Yam et al., 2004; Urtasun and Fua, 2004; Yamauchi et al., 2009; Zhao et al., 2006) With extremely excluding visual clues, these gait representations, which are formally parameterized as coordinates of human joints or customized vectors in most cases, are theoretically ‘clean’ against factors like carrying and dressing items. In recent literature, PoseGait (Liao et al., 2020) combines the 3D skeleton data with hand-crafted characteristics to overcome the viewpoint and clothing variations, GaitGraph (Teepe et al., 2021) introduces a graph convolution network for 2D skeleton-based gait representation learning, HMRGait (Li et al., 2020) fine-tunes a pre-trained human mesh recovery network to construct an end-to-end SMPL-based model. Despite the advances achieved on indoor OU-MVLP(Takemura et al., 2018), previous model-based methods still have not exhibited competitive performance on real-world gait datasets.
Appearance-based gait recognition methods mostly learn gait features from silhouette or RGB images, leveraging informative visual characteristics.(Lin et al., 2021; Zhang et al., 2021, 2019; Han and Bhanu, 2005; Shiraga et al., 2016; Wu et al., 2016; Huang et al., 2021) With the advent of deep learning, current appearance-based approaches primarily concentrate on spatial feature extraction and gait temporal modeling. Specifically, GaitSet (Chao et al., 2021) innovatively treats the gait sequence as a set and employs a maximum function to compress the sequence of frame-level spatial features. Due to its simplicity and effectiveness, GaitSet has emerged as one of the most influential gait recognition works in recent years. GaitPart (Fan et al., 2020) meticulously explores the local details of input silhouettes and models temporal dependencies using the Micro-motion Capture Module. GaitGL (Lin et al., 2022) argues that spatially global gait representations often overlook important details, while local region-based descriptors fail to capture relationships among neighboring parts. Consequently, GaitGL introduces global and local convolution layers. More recently, DeepGaitV2 (Fan et al., 2023a) presents a unified perspective to explore how to construct deep models for outdoor gait recognition, bringing a breakthrough improvement on the challenging Gait3D and GREW.
@@ 2.2. Gait Recognition Methods
Additionally, there are also some progressive multi-modal gait recognition methods aiming to fuse multiple modalities to create a more comprehensive representation of gait. Specifically, SMPLGait (Zheng et al., 2022) exploited the 3D geometrical information from the SMPL model to enhance the gait appearance feature learning. BiFusion(Peng et al., 2024) integrated silhouettes and skeletons, using a graph convolutional network (GCN) with a multi-scale gait graph (MSGG) to extract complementary features from both modalities. Skeletongait++(Fan et al., 2024) employed the Skeleton Map to represent the skeleton and integrated it with the silhouette using an attention mechanism.
Figure 2. Overview of the PSGait framework. During data preprocessing, the silhouettes and skeletons are generated from video frames. In multimodal fusion, after the parsing section, the Parsing Skeletons are fused with silhouettes. The fused data is then fed into the gait recognition model for individual differentiation. The parsing section refines body part representations through selected points, creating circular and linear representations to enhance the robustness and accuracy of gait recognition in the wild. Both the data preprocessing and gait recognition models are interchangeable.
2.3. Human parsing in gait recognition
Human parsing is a specific task of semantic segmentation for pixel-level classification of human parts(Liu et al., 2014; Luo et al., 2018; Zhou et al., 2018; Wang et al., 2020; Liu et al., 2022; Gong et al., 2018, 2017). Human parsing in gait recognition involves segmenting a person’s body into distinct regions or parts and using these parts to enhance gait recognition accuracy. This fine-grained approach enables the capture of body part-specific motion patterns, which are critical for recognizing individuals under challenging conditions, such as occlusions, clothing changes, or varying walking styles. Human parsing techniques typically rely on semantic segmentation or human body pose estimation, which involves detecting and labeling body parts in visual representations (e.g., RGB images or silhouettes).
@@ 2.3. Human parsing in gait recognition
Recent research in this domain has focused on leveraging both global and local features to improve gait recognition. ParsingGait (Zheng et al., 2023) combines convolutional neural networks (CNNs) and graph convolutional networks (GCNs) to capture both global and local gait features, thereby improving recognition performance under complex real-world conditions. Similarly, GaitParsing (Wang et al., 2023a) decomposes gait sequences into specific body parts using human semantic parsing. This approach employs a dual-branch feature extraction network to process both holistic and part-level features, allowing it to better capture fine-grained motion patterns. LandmarkGait (Wang et al., 2023b) introduces an unsupervised method that converts silhouettes into consistent landmarks, which can then be used for fine-grained parsing. This approach also incorporates a multi-scale feature extractor to capture both global and part-level features, further enhancing the gait recognition process.
3. METHOD
@@
3.1. Parsing Skeleton Representation
In this section, we introduce the definition of the novel representation, i.e., Parsing Skeleton, for gait recognition. Given a human walking sequence of frames $S=\{\mathbf{C}^{i}\}^{N}_{i=1}$, where $\mathbf{C}^{i}\in\mathbb{R}^{K\times 2}$ is one coordinates frame and $N$ is the length of the sequence, we can feed each frame $\mathbf{C}^{i}$ into the parsing Section $P(\cdot)$ to obtain the Parsing Skeleton frame $\textbf{x}^{i}=P(\textbf{C}^{i})$. The parsing section $P(\cdot)$ can set $K$ body parts (including the background) with different colors for each pixel in an image of the same size as $\textbf{C}^{i}$. Therefore, the Parsing Skeleton sequence can be formulated as:
$X=\{\mathbf{x}^{i}\}^{N}_{i=1},$
where $\textbf{x}^{i}\in\mathbb{R}^{H\times W}$ is the Parsing Skeleton frame. For each pixel $p^{j}$ of $\mathbf{x}^{i}$, we have $p^{j}\in\{0,1,...,K-1\}$ where 00 denotes the background.
Based on the above definition, we compare the proposed Parsing Skeleton representation with the binary silhouette from the viewpoint of information theory. Inspired by the definition of information entropy proposed by Claude E. Shannon (Shannon, 1948), the pixel-level information entropy of the Parsing Skeleton can be denoted as
$\mathcal{H}=-\sum_{k=0}^{K}p_{k}\cdot log(p_{k}),$
where pk is the possibility of one pixel belonging to the k-th class. For the conventional binary silhouette, each pixel only belongs to the foreground or the background, i.e., the $K=2$, so the information of the binary pixel is only one bit. The combination of Parsing Skeleton with silhouette, where K = 13, leads to a significantly higher information entropy compared to other gait recognition methods. This is due to the richer representation of 13 distinct human parts as shown in Fig. 2, allowing for a more complete capture of gait information and enhancing the model’s ability to distinguish subtle differences in gait patterns. Therefore, we consider Parsing Skeleton as a gait representation with higher information entropy to encode the fine-grained dynamics of body parts during walking.
3.2. Parsing Section
@@
$\text{line}((x_{1},y_{1}),(x_{2},y_{2}),w,c)=\left\{(u,v)\mid d\leq\frac{w}{2}\right\}$
@@ 3.3. Fusion Strategy
$F_{\text{overlay}}(x,y)=\max\left(M(x,y),S(x,y)\right)$
@@
Figure 3. Visualization of different gait representations. (a) Raw RGB frames from a walking sequence. (b) Binary silhouette images extracted via background subtraction. (c) 2D skeleton keypoints estimated from the RGB frames. (d) Our proposed Parsing Skeleton, which integrates part-level semantics and structural connectivity, providing richer spatial and motion cues for gait recognition. (e) Parsing Skeleton combined with silhouettes — the final fused representation used in PSGait, combining part-level semantics and body contours to enhance discriminability and structural consistency in gait recognition.
@@
$F_{\text{multi}}(x,y,c)=\begin{cases}S_{c}(x,y),&\text{if }c=1,2,\dots,11\\ M(x,y),&\text{if }c=12\end{cases}$
@@
3.4. Data Augmentation
To improve generalization and reduce overfitting, we apply data augmentation during training. Since gait sequences are already spatially aligned, we adopt a general image augmentation strategy following (Fan et al., 2023b), including: (1)Horizontal Flipping. Simulates varying walking directions without altering gait features, enhancing robustness to orientation changes. (2)Affine and Perspective Transformations. Introduces minor viewpoint variations through random geometric distortions, improving model adaptability to different camera angles. (3)Random Erasure. Randomly occludes image regions to mimic occlusion and background noise, encouraging the model to rely on global rather than local features. These strategies diversify the training data and improve the model’s robustness across real-world conditions.
Table 1. Quantitative comparison of gait recognition methods conducted across two authoritative datasets, CCPG(Li et al., 2023b) and Gait3D(Zheng et al., 2022). The best performances are indicated in bold, and the second-best methods are underlined. †We modified Skeletongait++ to match the settings used in GaitBase.
Input Method Source Testing Datasets